OpenStudy (anonymous):

How do i Factor completely: x2 − 3x + 7

OpenStudy (anonymous):

Bro here discriminant (b^2-4ac) is negative we will not get real roots but the roots are imaginary dude

jhonyy9 (jhonyy9):

x_1,_2=(3+/- sqrt(9-28))/2 =(3+/- sqrt(-19))/2=(3+/-isqrt19)/2 - so than x2 -3x +7 =(2x-(3-isqrt19))(2x-(3+isqrt19))